The length of a rectangle exceeds its width by 6 units. If each dimension were increased by 3 units,the area would be increased by 57 square units. Find the dimensions of the rectangle.
Answer by fractalier(6550)

Area of a rectangle is A = length times width...A=LW...
Here L = W + 6...substituting we get
A=(W+6)W
Now, with the increase
(W+9)(W+3)=(W+6)W + 57
Solve this...
W^2 + 12W + 27 = W^2 + 6W + 57
6W = 30
W = 5 and then
L = 11